# Planetary Edge Trends (PET)

### v1

@@ 3.4 Effect of Observational Bias
For our observational study, a crucial step is to assess whether observational selection bias could potentially influence our sample analysis results. Our planet samples are from the $Kepler$ survey, the major observational bias stems from the transit geometric selection effect and the detection efficiency of the $Kepler$ telescope. Since the detection efficiency does not vary significantly with stellar temperature or mass (as shown in Fig. 11 of Yang et al., 2020), we believe it will not significantly affect the correlation between the inner edge and stellar mass. Thus, in this section, we focus on the effect of transit geometric selection bias on the correlation between the inner edge and stellar mass. Specifically, assuming that the inner edge and stellar mass are uncorrelated, we perform simulations to model the transit selection process and determine whether any correlation or lack thereof is induced.
The specific analysis steps are as follows, and the results are shown in Fig. 6:
  1. 1.
    A stellar mass ($M_{\star}$) and its corresponding stellar radius ($R_{\star}$) are randomly selected from the observational data, followed by a random selection of the inner edge ($a_{\text{in}}$). Unlike the observational data, we shuffle the pairing between $a_{\text{in}}$ and $M_{\star}$ by randomly selecting them independently from the observational data.
  2. 2.
    The normalized transit probability is calculated using the following formula:
    $P_{\text{transit}}=\frac{R_{\star}}{a_{\text{in}}}\Big{/}P_{\text{max}},$ (10)
    where:
    $P_{\text{max}}=\frac{R_{\star(\text{max})}}{a_{\text{in(min)}}}.$ (11)
    $P_{\text{max}}$ represents the maximum possible transit probability in the simulation. $R_{\star(\text{max})}$ represents the largest $R_{\star}$ in the sample, and $a_{\text{in(min)}}$ represents the smallest $a_{\text{in}}$ in the sample.
  3. 3.
    A random number is generated from a uniform distribution between (0, 1). If $P_{\text{transit}}$ is greater than the generated random number, then the generated data in Step 1 (i.e., $M_{\star}$, $R_{\star}$, and $a_{\text{in}}$) are retained. Otherwise, we return to Step 1 to regenerate another set of $M_{\star}$, $R_{\star}$, and $a_{\text{in}}$.
This process is continued until the same number of sample sets ($N_{\text{obs}}$) matching the observational data is obtained (points as shown in the bottom middle panel of Fig. 6). The $N_{\text{obs}}$ sets of ($M_{\star}$, $R_{\star}$, $a_{\text{in}}$) are fit to determine the correlation $\gamma_{1}$ power-law index between $M_{\star}$ and $a_{\text{in}}$ (orange dashed line as shown in the bottom middle panel of Fig. 6).
@@
According to the result in the bottom middle panel of Fig. 6, we present a typical case to demonstrate the effect of transit selection bias on the correlation analysis between the inner edge and stellar mass. We observe that transit observational selection bias has two effects. On the one hand, smaller stellar masses (thus smaller radii) and larger inner edges (represented by the dark points in the upper left of the bottom middle panel) have lower $P_{\text{transit}}$, causing transit observational selection bias to filter out these data points. This reduces the number of points in the upper left of the bottom middle panel and introduces a positive correlation between $M_{\star}$ and $a_{\text{in}}$. On the other hand, areas with high data density and high $P_{\text{transit}}$ (indicated by the bright points in the lower right of the bottom middle panel) are given greater weight. This effect counteracts the previous one and simultaneously weakens the correlation between $M_{\star}$ and $a_{\text{in}}$. Consequently, the correlation between $M_{\star}$ and $a_{\text{in}}$, induced by the transit observational selection bias, is minimal, with $\gamma_{1}$ approaching zero (0.003).
Table 2: Power-law Index of Result and Model
@@
$1/3\sim 0.333$ Mulders et al. (2015)
All Multiple Systems $0.812_{-0.078}^{+0.085}$ ($0.634_{-0.039}^{+0.039}$) This Work2
Rp $<$ Rgap (super-Earths) $0.566_{-0.108}^{+0.100}$ ($0.543_{-0.052}^{+0.053}$)
Rp $\geq$ Rgap (sub-Neptunes) $0.667_{-0.182}^{+0.173}$ ($0.384_{-0.088}^{+0.087}$)
Mixed Multiple Systems $1.115_{-0.072}^{+0.082}$ ($0.950_{-0.045}^{+0.046}$)
Theoretical Model Sublimation (active, ${\alpha}=2$)3 $11/9\sim 1.222$ Liu et al. (2019)
@@
Sublimation (active, ${\alpha}=1$)3 $7/9\sim 0.778$ Liu et al. (2019)
Stellar Tides $9/13\sim 0.692$ Jackson et al. (2009)
@@
Co-rotation Radius $1/3\sim 0.333$ Mulders et al. (2015)
Planetary Tides $3/13\sim 0.231$ Jackson et al. (2009)
@@
The entire process is repeated 1,000 times to produce the distribution of the correlation $\gamma_{1}$ power-law index for $M_{\star}$ and $a_{\text{in}}$ (illustrated in the bottom right panel of Fig. 6). As can be seen, the median power-law index of the correlation distribution is 0.003. In contrast, our observational results show a positive correlation between the inner edge and stellar mass of 0.634 (the correlation for all multiple systems, as shown in the top left panel of Fig. 3), which is significantly larger than the median of the distribution. Additionally, in 1,000 simulations, no positive correlation with $\gamma_{1}$ reaching 0.634 is produced. Therefore, observational selection bias does not affect our analysis of the relationship between stellar mass and the inner edge.
In order to check whether the above results rely on the observational data, we repeat the above test with a modification to the analysis in Step 1. We provide the flat results, where the data are generated uniformly from the plane of $M_{\star}-a_{\text{in}}$ (points as shown in the bottom left panel of Fig. 6), rather than from the observational sample. In this case, $M_{\star}$ and $a_{\text{in}}$ are randomly selected within their respective ranges, while $R_{\star}$ is recalculated based on the fit between mass and radius from the observational data. The result is shown in the bottom left panel of Fig. 6. We present a typical case to demonstrate the effect of transit selection bias on the correlation analysis between the inner edge and mass, where $\gamma_{1}$ is also 0.003 (black dashed line as shown in the bottom left panel of Fig. 6). The result is almost identical to the bottom middle panel, indicating that the test result is robust regardless of the specific data.
To ensure thorough verification, we repeat the above test but change the number of sample sets ($N_{\text{test}}$ = 10,000). As shown in the top left and middle panels of Fig. 6, the correlation $\gamma_{1}$ of the relationship between stellar mass and the inner edge is close to zero (black dashed line and orange dashed line). The distribution in the top right panel of Fig. 6 indicates that the median power-law index of the correlation distribution is 0.001, which is far from the observed positive correlation $\gamma_{1}$ (0.634). In addition, compared to the bottom right panel, the distribution of $\gamma_{1}$ in the top right panel is more concentrated due to the use of a larger number of sample sets (10,000) in the simulation.
To summarize the above tests, we find that the transit selection bias does not induce a significant correlation between stellar mass and the inner edge.
4 Discussion
Figure 7: The correlation $\gamma_{1}$ comparison between different theoretical models and observational result. The green band represents the result of the dataset’s All Multiple Systems population obtained through the MLR model (same as the top panel in Fig. 5, with points having the same meaning). The dotted and dotted-dashed lines in different colors represent the power-law indices under various theoretical models, arranged in descending order from top to bottom. For comparison of correlation $\gamma_{1}$, the intercepts of the models are standardized with the intercept of the observational result.
@@ 4.1 Implications for Planet Formation and Evolution
By conducting the above analysis on the dataset across four populations (All Multiple Systems, Super-Earth Systems, Sub-Neptune Systems, and Mixed Multiple Systems), we generally confirm the results of previous studies (Plavchan & Bilinski, 2013; Mulders et al., 2015) and find that the relationship between the inner edge and stellar mass remains qualitatively consistent, with the inner edge position increasing with stellar mass. However, quantitatively, the degree of dependence between the inner edge and stellar mass appears to vary across different populations. Furthermore, by isolating the effect of stellar mass through correcting for metallicity, the influence of stellar mass on the inner edge is further strengthened, showing a correlation index of $0.6-1.1$ with a power-law relationship (Figs. 3, 5, and Table 2). Additionally, observational selection bias does not impact our results. In the following, we discuss the implications of our results for planet formation and evolution.
@@
There are a number of theoretical models in the literature exploring the relationship between the inner edge and stellar mass. For the pre-main-sequence (PMS) dust sublimation radius of an actively accreting disk, as derived from Eq. (9) in Liu et al. (2019), the relationship is given by $a_{\text{in}}\propto M_{\star}^{(3+4\alpha)/9}$, where $\alpha$ represents the power-law index of the scaling between the mass accretion rate and stellar mass ($\dot{M}_{\text{acc}}\propto M_{\star}^{\alpha}$). Observations suggest $\alpha=2$ (Alcalá et al., 2014), and we also examine the case where $\alpha=1$. The corresponding power-law indices are $a_{\text{in}}\propto M_{\star}^{11/9}$ and $a_{\text{in}}\propto M_{\star}^{7/9}$, respectively. For the PMS dust sublimation radius of a passive disk, $a_{\text{in}}\propto L_{\star}^{1/2}$ (Dullemond et al., 2001; Tripathi et al., 2017) and $L_{\star}\propto M_{\star}^{k}$, where $k$ ranges from 1 to 2 during the PMS stars. At earlier stages ($t<1\,\text{Myr}$), $k$ is approximately 2, gradually declines toward 1 by $t\sim 10\,\text{Myr}$. Thus, we consider two values for $k$: 1 and 2. In that case, $a_{\text{in}}\propto M_{\star}^{0.5}$ for $k=1$, and $a_{\text{in}}\propto M_{\star}^{1.0}$ for $k=2$, based on the temperature profile of the minimum mass solar nebula (MMSN). For planet destruction by stellar tides and by planetary tides maintained through secular interactions, the scaling follows Eqs. (15) and (16) in Mulders et al. (2015), which are derived from Eq. (1) in Jackson et al. (2009). Assuming $R_{\star}\propto M_{\star}$, the respective power-law indices are $a_{\text{in}}\propto M_{\star}^{9/13}$ and $a_{\text{in}}\propto M_{\star}^{3/13}$. The PMS co-rotation radius represents the location where the inner edge truncates at a frequency of planetary revolution synchronized with the stellar rotation frequency (Lee & Chiang, 2017). In a state of Keplerian rotation, a relationship between stellar mass and the inner edge position is obtained as $a_{\text{in}}\propto M_{\star}^{1/3}$ (Mulders et al., 2015).
Figure 8: A schematic comparison of the conclusions of this work and Mendigutía et al. (2024), revealing that the inner edges of different planetary systems correspond to different mechanisms. The results obtained in this work for small planets align more closely with the dust sublimation model, whereas their results obtained for giant planets are more consistent with migration up to the gas-truncation radius.
@@
Table 2 summarizes these theoretical models, along with the observational results from previous work and this study. Fig. 7 illustrates the comparison between observational result and the theoretical models described above, using the same intercept (here, the intercept obtained from the MLR model for all multiple systems, $\log(\gamma_{0})=-1.195$ in Eq. (8)). In Fig. 7, we observe that the Sublimation (active, ${\mathrm{\alpha}}=1$) model ($\sim 0.778$) provides the best fit, matching the observed correlation slope ($0.812_{-0.078}^{+0.085}$) within one sigma consistency at the same intercept. Additionally, the Stellar Tides model is also close to the observed result. However, the Sublimation (active, ${\mathrm{\alpha}}=2$) model and the Sublimation (passive, $k=2$) model show larger slopes, while the Sublimation (passive, $k=1$) model, the Co-rotation Radius model, and the Planetary Tides model show smaller slopes.
@@
In summary, dust sublimation is likely a key factor shaping the dependence of the inner edge of small planets on stellar mass. In addition, for the $\gamma_{1}$ values across all populations, which fall within the range of approximately $0.6-1.1$, we propose that both dust sublimation and stellar tides, having similar effects, could play a role in shaping this dependence. If these two mechanisms simultaneously influence the inner edge, their combined action might result in a stronger correlation. In future work, more theoretical models will be needed to explore the outcomes of these two mechanisms acting together.

### v2

@@ 3.4 Effect of observational bias
For our observational study, a crucial step was to assess whether observational selection bias could potentially influence our sample analysis results. Our planet samples are from the $Kepler$ survey, the major observational bias stems from the transit geometric selection effect and the detection efficiency of the $Kepler$ telescope. Since the detection efficiency does not vary significantly with stellar temperature or mass (as shown in Fig. 11 of Yang et al., 2020), we believe it will not significantly affect the correlation between the inner edge and stellar mass. Thus, in this section, we focus on the effect of transit geometric selection bias on the correlation between the inner edge and stellar mass. Specifically, assuming that the inner edge and stellar mass are uncorrelated, we performed simulations to model the transit selection process and determined whether any correlation or lack thereof is induced.
The specific analysis steps are as follows, and the results are shown in Fig. 6. (i) A stellar mass ($M_{\star}$) and its corresponding stellar radius ($R_{\star}$) were randomly selected from the observational data, followed by a random selection of the inner edge ($a_{\text{in}}$). Unlike the observational data, we shuffled the pairing between $a_{\text{in}}$ and $M_{\star}$ by randomly selecting them independently from the observational data. (ii) The normalized transit probability was calculated using the following formula:
$P_{\text{transit}}=\frac{R_{\star}}{a_{\text{in}}}\Big{/}P_{\text{max}},$ (10)
where
$P_{\text{max}}=\frac{R_{\star(\text{max})}}{a_{\text{in(min)}}}.$ (11)
Here, $P_{\text{max}}$ represents the maximum possible transit probability in the simulation, $R_{\star(\text{max})}$ represents the largest $R_{\star}$ in the sample, and $a_{\text{in(min)}}$ represents the smallest $a_{\text{in}}$ in the sample. (iii) A random number was generated from a uniform distribution between zero and one. If $P_{\text{transit}}$ was greater than the random number, then the data generated in the first step (i.e., $M_{\star}$, $R_{\star}$, and $a_{\text{in}}$) were retained. Otherwise, we returned to the first step to regenerate another set of $M_{\star}$, $R_{\star}$, and $a_{\text{in}}$. This is the core step of this section. We simulated observations based on the relative probability of transits, assigning different weights to individual data points, where the weights reflected their relative probabilities. Targets with higher probabilities—meaning they are more likely to be observed—were selected for the next stage of analysis, which focused on observational selection bias.
This process was continued until the same number of sample sets ($N_{\text{obs}}$) matching the observational data was obtained (points as shown in the bottom-middle panel of Fig. 6). The $N_{\text{obs}}$ sets of ($M_{\star}$, $R_{\star}$, $a_{\text{in}}$) were fit to determine the correlation $\gamma_{1}$ power-law index between $M_{\star}$ and $a_{\text{in}}$ (orange dashed line shown in the bottom-middle panel of Fig. 6).
@@
According to the result in the bottom-middle panel of Fig. 6, we present a typical case to demonstrate the effect of transit selection bias on the correlation analysis between the inner edge and stellar mass. We observe that transit observational selection bias has two effects. On the one hand, smaller stellar masses (thus smaller radii) and larger inner edges (represented by the dark points in the upper-left of the bottom-middle panel) have lower $P_{\text{transit}}$, causing transit observational selection bias to filter out these data points. This reduces the number of points in the upper-left of the bottom-middle panel and introduces a positive correlation between $M_{\star}$ and $a_{\text{in}}$. On the other hand, regions with high data density and high $P_{\text{transit}}$ (indicated by the bright points in the lower-right of the bottom-middle panel) carry greater weight in the data fitting process. Without the effect from the lower-right corner, the fitted slope would be steeper due to observational selection bias. It is precisely this effect that leads to a flatter fitted line. This effect counteracts the previous one and simultaneously weakens the correlation between $M_{\star}$ and $a_{\text{in}}$. Consequently, the correlation between $M_{\star}$ and $a_{\text{in}}$ induced by the transit observational selection bias is minimal ($\gamma_{1}=0.001$).
The entire process was repeated 1,000 times to produce the distribution of the correlation $\gamma_{1}$ power-law index for $M_{\star}$ and $a_{\text{in}}$ (illustrated in the bottom-right panel of Fig. 6). As can be seen, the median power-law index of the correlation distribution is 0.001. In contrast, our observational results show a positive correlation between the inner edge and stellar mass of 0.63 (the correlation for all multiple systems, as shown in the top-left panel of Fig. 3), which is significantly larger than the median of the distribution. Additionally, in 1,000 simulations, almost no positive correlation with $\gamma_{1}$ reaching 0.63 is produced. Therefore, we conclude that while observational selection bias may influence the slope, it is not solely responsible for the observed correlation.
In order to check whether the above results rely on the observational data, we repeated the above test with a modification to the analysis in the first step. We provided the flat results, where the data were generated uniformly from the plane of $M_{\star}-a_{\text{in}}$ (points as shown in the bottom-left panel of Fig. 6), rather than from the observational sample. In this case, $M_{\star}$ and $a_{\text{in}}$ are randomly selected within their respective ranges, while $R_{\star}$ is recalculated based on the fit between mass and radius from the observational data. The result is shown in the bottom-left panel of Fig. 6. We present a typical case to demonstrate the effect of transit selection bias on the correlation analysis between the inner edge and mass, where $\gamma_{1}$ is $0.002$ (black dashed line as shown in the bottom-left panel of Fig. 6). The result is almost identical to the bottom-middle panel, indicating that the test result is robust regardless of the specific data.
To ensure a thorough verification, we repeated the above test but changed the number of sample sets ($N_{\text{test}}$ = 10,000). As shown in the top-left and middle panels of Fig. 6, the correlation $\gamma_{1}$ of the relationship between stellar mass and the inner edge is close to zero (black dashed line and orange dashed line). The distribution in the top-right panel of Fig. 6 indicates that the median power-law index of the correlation distribution is 0.001, which is far from the observed positive correlation $\gamma_{1}$ (0.63). In addition, compared to the bottom-right panel, the distribution of $\gamma_{1}$ in the top-right panel is more concentrated due to the use of a larger number of sample sets (10,000) in the simulation.
To summarize the above tests, we find that the transit selection bias does not induce a significant correlation. This suggests that the observed correlation between stellar mass and the inner edge is not driven by observational bias.
4 Discussion
Table 2: Power-law index of the result and model.
@@ 4 Discussion
$1/3\sim{\color[rgb]{0,0,0}0.33}$ Mulders et al. (2015)
all multiple systems $0.81_{-0.08}^{+0.09}$ ($0.63_{-0.04}^{+0.04}$) This Work2
Rp $<$ Rgap (super-Earths) $0.57_{-0.11}^{+0.10}$ ($0.54_{-0.05}^{+0.05}$)
Rp $\geq$ Rgap (sub-Neptunes) $0.67_{-0.18}^{+0.17}$ ($0.38_{-0.09}^{+0.09}$)
mixed multiple systems $1.12_{-0.07}^{+0.08}$ ($0.95_{-0.05}^{+0.05}$)
Theoretical Model Sublimation (active, ${\alpha}=2$)3 $11/9\sim{\color[rgb]{0,0,0}1.22}$ Liu et al. (2019)
@@
Sublimation (active, ${\alpha}=1$)3 $7/9\sim{\color[rgb]{0,0,0}0.78}$ Liu et al. (2019)
Stellar Tides $9/13\sim{\color[rgb]{0,0,0}0.69}$ Jackson et al. (2009)
@@
Co-rotation Radius $1/3\sim{\color[rgb]{0,0,0}0.33}$ Mulders et al. (2015)
Planetary Tides $3/13\sim{\color[rgb]{0,0,0}0.23}$ Jackson et al. (2009)
@@
Figure 7: Correlation $\gamma_{1}$ comparison between different theoretical models and the observational result. The green band represents the result of the dataset’s all multiple systems population obtained through the MLR model (same as the top panel in Fig. 5, with points having the same meaning). The dotted and dotted-dashed lines in different colors represent the power-law indices under various theoretical models, arranged in descending order from top to bottom. For comparison of correlation $\gamma_{1}$, the intercepts of the models are standardized with the intercept of the observational result.
@@ 4.1 Implications for planet formation and evolution
By conducting the above analysis on the dataset across four populations (all multiple systems, super-Earth systems, sub-Neptune systems, and mixed multiple systems), we generally confirm the results of previous studies (Plavchan & Bilinski, 2013; Mulders et al., 2015) and find that the relationship between the inner edge and stellar mass remains qualitatively consistent, with the inner edge position increasing with stellar mass. However, quantitatively, the degree of dependence between the inner edge and stellar mass appears to vary across different populations. Furthermore, by isolating the effect of stellar mass through correcting for metallicity, the influence of stellar mass on the inner edge is further strengthened, showing a correlation index ($\gamma_{1}$) of $0.6-1.1$ with a power-law relationship (Figs. 3, 5, and Table 2). Additionally, observational selection bias does not impact our results. In the following, we discuss the implications of our results for planet formation and evolution.
@@
There are a number of theoretical models in the literature exploring the relationship between the inner edge and stellar mass. For the pre-main-sequence dust sublimation radius of an actively accreting disk, as derived from Eq. (9) in Liu et al. (2019), the relationship is given by $a_{\text{in}}\propto M_{\star}^{(3+4\alpha)/9}$, where $\alpha$ represents the power-law index of the scaling between the mass accretion rate and stellar mass ($\dot{M}_{\text{acc}}\propto M_{\star}^{\alpha}$). Observations suggest $\alpha=2$ (Alcalá et al., 2014), and we also examined the case where $\alpha=1$. The corresponding power-law indices are $a_{\text{in}}\propto M_{\star}^{11/9}$ and $a_{\text{in}}\propto M_{\star}^{7/9}$, respectively. For the pre-main-sequence dust sublimation radius of a passive disk, $a_{\text{in}}\propto L_{\star}^{1/2}$ (Dullemond et al., 2001; Tripathi et al., 2017) and $L_{\star}\propto M_{\star}^{k}$, where $k$ ranges from one to two during the pre-main-sequence stars. At earlier stages ($t<1\,\text{Myr}$), $k$ is approximately two, gradually declines toward one by $t\sim 10\,\text{Myr}$. Thus, we considered two values for $k$: one and two. In that case, $a_{\text{in}}\propto M_{\star}^{0.5}$ for $k=1$, and $a_{\text{in}}\propto M_{\star}^{1.0}$ for $k=2$, based on the temperature profile of the minimum mass solar nebula. For planet destruction by stellar tides and by planetary tides maintained through secular interactions, the scaling follows Eqs. (15) and (16) in Mulders et al. (2015), which are derived from Eq. (1) in Jackson et al. (2009). Assuming $R_{\star}\propto M_{\star}$, the respective power-law indices are $a_{\text{in}}\propto M_{\star}^{9/13}$ and $a_{\text{in}}\propto M_{\star}^{3/13}$. The pre-main-sequence co-rotation radius represents the location where the inner edge truncates at a frequency of planetary revolution synchronized with the stellar rotation frequency (Privitera et al., 2016; Lee & Chiang, 2017). In a state of Keplerian rotation, a relationship between stellar mass and the inner edge position is obtained as $a_{\text{in}}\propto M_{\star}^{1/3}$ (Mulders et al., 2015).
@@
Table 2 summarizes these theoretical models, along with the observational results from previous work and this study. Figure 7 illustrates the comparison between observational result and the theoretical models described above, using the same intercept (here, the intercept obtained from the MLR model for all multiple systems, $\log(\gamma_{0,amf})=-1.195$ in Eq. (8)). In Fig. 7, we observe that the Sublimation (active, ${\mathrm{\alpha}}=1$) model ($\sim{\color[rgb]{0,0,0}0.78}$) provides the best fit, matching the observed correlation slope ($0.81_{-0.08}^{+0.09}$) within one sigma consistency at the same intercept. Additionally, the Stellar Tides model is also close to the observed result. However, the Sublimation (active, ${\mathrm{\alpha}}=2$) model and the Sublimation (passive, $k=2$) model show larger slopes, while the Sublimation (passive, $k=1$) model, the Co-rotation Radius model, and the Planetary Tides model show smaller slopes.
@@
In summary, dust sublimation is likely a key factor shaping the dependence of the inner edge of small planets on stellar mass. In addition, for the $\gamma_{1}$ values across all populations, which fall within the range of approximately $0.6-1.1$, we propose that both dust sublimation and stellar tides, having similar effects, could play a role in shaping this dependence. If these two mechanisms simultaneously influence the inner edge, their combined action might result in a stronger correlation. In future work, more theoretical models will be needed to explore the outcomes of these two mechanisms acting together. Given that existing models can provide quantitative relationships between stellar mass and the inner edge during the pre-main-sequence phase, this study focuses on theoretical models related to planet formation. However, we emphasize that the evolution of planetary systems can also play a significant role. We believe that the observational results likely reflect the combined effects of both planetary system formation and evolution (see Sect. 4.3). Therefore, in the models discussed, tidal effects—such as stellar tides and planetary tides (i.e., evolutionary models)—should also be considered. However, it is important to note that this process primarily contributes to the formation of ultra-short-period planets, which have already been excluded from our study (see Sect. 2.2).